Please find the following limit
lim(x->inf) (x+2/x+1)x

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
L = \lim_{x\rightarrow \infty }(\frac{x+2}{x+1})^{x}
L = \lim_{x\rightarrow \infty }(1+\frac{1}{x+1})^{x}
L = \lim_{x\rightarrow \infty }(1+\frac{1}{x+1})^{\frac{x+1}{x+1}.x}
L = \lim_{x\rightarrow \infty }(1+\frac{1}{x+1})^{\frac{1}{x+1}} = e
L = e^{\lim_{x\rightarrow \infty }{\frac{x}{x+1}}}
L = e^1 = e